Suppose y varies jointly as x and z and y=36, when x=1.2 and z=2. Find y when x=0.4 and z=3. How would you solve this problem ?
Answer by John10(297)

The formula is y = k(xz) where k is the variance constant
Now we find k = ?
k = y/(xz) = 36/(1.2* 2) = 15
Now use the new k to find y:
y = 15(0.4*3) = 18
John10:)